I'm stumped. Solve for Y
4(2y-5)=2(4y-10)
8y-20=8y-20
8y-20+20=8y-20+20
8y=8y
1/8*8y=1/8*8y
y=y
Were am I going wrong?

Answer by venugopalramana(3286)

THERE IS NOTHING WRONG WITH YOUR ANSWER..THE PROBLEM GIVEN IS LIKE THAT.. IT IS AN IDENTITY...WHEN LEFT HAND SIDE OF AN EQUATION IS SAME AS THAT ON THE RIGHT HAND SIDE IT IS CALLED AN IDENTITY..IT IS TRUE FOR ANY VALUE OF THE VARIABLE INVOLVED.SO YOU CAN SAY THAT Y = 0,1,-1,1/2....ETC..ANY AND EVERY VALUE. SOME MORE EXAMPLES OF IDENTITIES ARE
(X+Y)^2=X^2+Y^2+2*X*Y WHICH IS TRUE FOR ANY VALUE OF X AND Y.